OpenStudy (anonymous):

You have already invested $400 in a stock with an annual return of 11%. How much of an additional $1,200 should be invested at 12% and how much at 6% so that the total return on the entire $1,600 is 9%? (Round each answer to the nearest cent.)

OpenStudy (anonymous):

Please help guys

OpenStudy (anonymous):

The annual return on $400 in stock is 0.11*400 = $44

OpenStudy (anonymous):

the total return you want is 0.09 * 1600 = $144

OpenStudy (anonymous):

let a be the amount invested at 12% and b be the amount invested at 6% 0.12a + 0.06b = 144 - 44

OpenStudy (anonymous):

12a + 6b = 10000

OpenStudy (anonymous):

you mean 100?

OpenStudy (anonymous):

a + b = 1200

OpenStudy (anonymous):

put those 2 equations together and solve for a and b: 12a + 6b = 10000 a + b = 1200

OpenStudy (anonymous):

Using elimination, I got a = $466.67 and b = $733.33

OpenStudy (anonymous):

So, out of an additional $1200, invest $466.67 at a rate of 12% and $733.33 at a rate of 6%

OpenStudy (anonymous):

Thats correct. Thanks!

OpenStudy (anonymous):

Fill in this table.